OpenStudy (anonymous):

find the greatest common factor for the pair of numbers 105,175

OpenStudy (anonymous):

3*5*7=105 5*5*7 = 175 we are looking for the ones that are common so 5 and 7 5*7 = 35 which is youre largest common factor